OpenStudy (i_love_my_nieces):

FAN AND MEDAL!!!!!!!!!!!!!!!!!!!! You are having a discussion about sequences with your classmates. She insists that she the sequence 2, 3, 5, 8, 12 must be either arithmetic or geometric. IS she correct of incorrect? Explain.

OpenStudy (i_love_my_nieces):

@mathmale

OpenStudy (cwrw238):

an arithmetic sequencs has terms which have a common difference eg 2, 4 , 6, 8 - each term is formed by adding 2 to the previous term another example would be 2,5,8,11 (common difference = 3) a geometric sequence has a common ratio - each term is formed from previous by multiplying by this ratio eg 2 , 4,8 ,16 common ratio is 2 so what do you think ?

OpenStudy (cwrw238):

is there a common difference for 2 3 5 8 12?

OpenStudy (i_love_my_nieces):

No, it have to answer the question

OpenStudy (cwrw238):

correct 3-2 = 1 5-3 = 2 8-5 = 3 the difference is increasing for each term so it is not arithmetic now try dividing second term by the first term and third term by the second to see if there is a common ratio

OpenStudy (cwrw238):

3 / 2 = 5 / 3 = 8 / 5 =

OpenStudy (cwrw238):

do these give the same result?

OpenStudy (cwrw238):

If not your answer is neither

OpenStudy (cwrw238):

So she is incorrect

OpenStudy (i_love_my_nieces):

I think she si corrent

OpenStudy (cwrw238):

no - try working out those divisions and see if they give the same result - I think you'll find they dont - so it cant be a geometric sequence either.

OpenStudy (cwrw238):

3/2 = 1.5 5/3 = 1.67